OpenStudy (anonymous):

Create a graph that represents the function f (x − 3), if f (x) = x^2 @Mertsj

OpenStudy (whpalmer4):

Do you know what the graph of \(y = f(x) = x^2\) looks like?

OpenStudy (whpalmer4):

This will be a translation of that graph in one of four directions: up, down, right, or left.

OpenStudy (whpalmer4):

Which one? If we add a number to the result of \(f(x)\), such as \(y = f(x)+1\), we make each point appear 1 unit higher on the graph, right? If we subtract a number from the result of \(f(x)\), such as \(y = f(x)-1\), we make each point appear 1 unit lower on the graph. Doesn't matter what the definition of \(f(x)\) actually is; we shift the graph up or down by the amount added or subtracted from the output of \(f(x)\).

OpenStudy (anonymous):

Uhh the second one?

OpenStudy (whpalmer4):

You could think of it as plotting the graph of \(y = f(x)\) on a sheet of clear plastic, and sliding it up and down over your graph paper. Now, that's the easier pair of translations to understand. Let's make a little table: x x^2 x^2+1 0 0 1 1 1 2 2 4 5 3 9 10 4 16 17 if you plot those points on a graph, drawing a separate curve through the x^2+1 points, you'll see that adding 1 to x^2 just reproduces the same curve, shifted up 1 unit.

OpenStudy (whpalmer4):

Blue line is x^2, purple line is x^2+1

OpenStudy (anonymous):

Oh wait, it gives me options.. Should I place the options?

OpenStudy (whpalmer4):

Now let's tackle the other case: we add or subtract to/from x before squaring it x x+1 x^2 (x+1)^2 0 1 0 1 1 2 1 4 2 3 4 9 3 4 9 16 4 5 16 25 Do you see what is happening? the (x+1)^2 column has the same values as the x^2 column, except where the x^2 column has a given value at x, the (x+1)^2 column has it at x-1!

OpenStudy (whpalmer4):

Nah, we don't need no steekin' options, we're going to learn how to do this all by ourselves!

OpenStudy (anonymous):

First one A, second one B, third one C, fourth one D.

OpenStudy (whpalmer4):

Here's a plot of those two curves: Again, blue is x^2, purple is (x+1)^2

OpenStudy (anonymous):

Ok.

OpenStudy (whpalmer4):

So when we add a number to the argument of the function, we shift the curve to the left. What do you think happens when we subtract a number from the argument (such as occurs in your problem)?

OpenStudy (anonymous):

Shifts the curve to the right?

OpenStudy (whpalmer4):

That's correct!

OpenStudy (whpalmer4):

So, if our original function is \[y=f(x) = x^2\] and we change it to \[y_{new} = f(x-3) = (x-3)^2\]the new graph looks like?

OpenStudy (anonymous):

Idk how to draw it lol

OpenStudy (whpalmer4):

Well, describe what happens to it...

OpenStudy (whpalmer4):

if you don't know what \(y = x^2\) looks like, make a little table as I did: x y=x^2 -4 16 -3 9 -2 4 -1 1 0 0 1 1 2 4 3 9 4 16 and plot those points and connect them with a smooth curve.

OpenStudy (whpalmer4):

then apply the change that subtracting 3 from the argument does (shifting it right or left) to figure out which of the answer choices is correct. Or, make a table that computes (x-3)^2 directly and plot those points: x y = (x-3)^2 ... 0 9 1 4 etc.

OpenStudy (anonymous):

Call me stupid but uhm, I'm confused. So bad. This is TOO advanced for me xD.

OpenStudy (whpalmer4):

no, review what we said and did: adding a number to the result \((y=f(x)+1)\) shifted the graph up. subtracting a number from the result \((y = f(x)-1)\) shifted the graph down. we aren't doing that here, so if any of the answer graphs showed the parabola shifted up or down, they would not be candidates to be a correct answer. however, it doesn't look like we have any of those, so we haven't accomplished any narrowing of the choices yet :-) adding a number to the argument of the function \((y=f(x+1))\) shifted the graph to the left. You can think of it this way: if there's some spot on the original graph that happens at \(x = 0\), the same spot has to be at \(x+1=0\) or \(x = -1\) on the shifted graph. subtracting a number from the argument of the function \((y = f(x-1))\) shifts the graph to the right. Again, something that happens at \(x = 0\) on the original now has to be at \(x-1 = 0\) or \(x = 1\).

OpenStudy (whpalmer4):

So, our function is \(y = f(x) = x^2\). We change it to \(y = f(x-3) = (x-3)^2\). That shifts the graph 3 units to the right. All you have to do is figure out which of those graphs represents the graph of \(y = x^2\), shifted 3 units to the right.

OpenStudy (whpalmer4):

You can do it! Unfortunately, I have to leave now to take my cat to the vet :-(